# Batched Nonparametric Contextual Bandits

### v1

@@ 5 Regret analysis for BaSEDB
Our proof of the regret upper bound is inspired by the framework developed in [39]. Our setting presents additional technical difficulty due to the batch constraint.
@@ 5.1 Two clean events
$\bar{\mathcal{I}_{C}}$ =$=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{1}|C|^{\beta}\right\},$
@@
In words, Lemma 3 guarantees that $\mathcal{A}_{C}$ happens with high probability if $E$ holds and $\mathcal{A}_{C^{\prime}}$ holds for all the ancestors $C\text{\textquoteright}$ ’ of $C$. See Section 5.5.2 for the proof.
5.2 Regret decomposition
@@
5.3 Controlling three terms
@@
5.3.1 Controlling $V_{i}$
Fix some $11\leq i\leq M-$ 1, and some bin $C\in\mathcal{B}_{i}$. On the event $\mathcal{G}_{C}$ we have $\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}\subseteq\bar{I}_{\mathsf{p}(C)}$, that is, for any $k\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$,
@@ 5.3.3 Last Batch
5.4 Putting things together
@@
5.5 Proofs for the clean events
@@ 5.5.3 Auxiliary lemmas
6 Proof of suboptimality of static binning
@@ 6.1 A helper lemma
6.2 Three failure cases for $g$
@@
where we have used the choice of $z$. So Theorem (3) holds with $\kappa=\epsilon$.
@@
So Theorem (3) holds with $\kappa=\epsilon/4$.
@@
We see that Theorem (3) holds with $\kappa=1/38$.
7 Discussion
Extensions to multiple arms.
With slight modification, our algorithm works for nonparametric contextual bandits with more than two arms. However, it remains unclear what the fundamental limits of batch learning are in this multi-armed case.
Improving the log factor.
Adapting to smoothness and margin parameters.
In practice, we do not always know the smoothness and the margin parameters. Can one develop a batch learning algorithm that can adapt to these unknown parameters? This question is intriguing because in the fully online setting, a similar adaptively binned successive elimination algorithm was proposed in [39] with the sole purpose of adapting to the margin parameter. [^3] We emphasize again that if adaptivity to the margin parameter is not needed, then static binning suffices for the online setting.
Acknowledgements

### v2

@@ Appendix B Proof of Theorem 3
Our proof of Theorem 3 is inspired by the framework developed in [39]. Our setting presents additional technical difficulty due to the batch constraint.
@@ B.1 Two clean events
¯$\bar{\mathcal{I}}_{C}$ =$=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{1}|C|^{\beta}\right\},$
@@
In words, Lemma 8 guarantees that $\mathcal{A}_{C}$ happens with high probability if $E$ holds and $\mathcal{A}_{C^{\prime}}$ holds for all the ancestors $C\text{\textquoteright}$ ’ of $C$. See Section B.5.2 for the proof.
B.2 Regret decomposition
@@
B.3 Controlling three terms
@@
B.3.1 Controlling $V_{i}$
Fix some $11\leq i\leq M-$ 1, and some bin $C\in\mathcal{B}_{i}$. On the event $\mathcal{G}_{C}$ we have $\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}\subseteq\bar{\mathcal{I}}_{\mathsf{p}(C)}$, that is, for any $k\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$,
@@ B.3.3 Last Batch
B.4 Putting things together
@@
B.5 Proofs for the clean events
@@ B.6 Auxiliary lemmas
Appendix C Proof of Theorem 4
@@ C.1 A helper lemma
C.2 Three failure cases for $g$
@@
where we have used the choice of $z$. So Theorem 4 holds with $\kappa=\epsilon$.
@@
So Theorem 4 holds with $\kappa=\epsilon/4$.
@@
We see that Theorem 4 holds with $\kappa=1/38$.
Appendix D Proof of Theorem 5
When $\beta=d=1$, we denote $\gamma(\alpha^{\star})=(\alpha^{\star}+1)/3$. Fix some small constant $\epsilon>0$ to be specified later. We deal with $M=2$ and $M=3$ separately. In both cases, we use the fact that the algorithm needs to provide its first batch size $t_{1}$ prior to the game, and design $\alpha^{\star}$ such that any choice of $t_{1}$ would fail.
D.1 When $M=2$
Under $M=2$, the theoretical optimal rate is $T^{(1-\gamma(\alpha^{\star}))/(1-\gamma(\alpha^{\star})^{2}})=T^{3/(\alpha^{\star}+4)}$.
Case of $t_{1}=T^{3/5+\epsilon}$.
Take $\alpha^{\star}=1$. Since fixed grid and adaptive grid are the same when $M=2$, by relation (9), we have
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\gtrsim\max\left\{t_{1},\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}\right\}\geq t_{1}=T^{\frac{3}{5}+\kappa_{1}},$
where $\kappa_{1}=\epsilon$.
Case of $t_{1}=T^{3/4-\epsilon}$.
Take $\alpha^{\star}=o(1)$. By relation (9), we have
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\gtrsim\max\left\{t_{1},\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}\right\}\geq\frac{T}{t_{1}^{\frac{\alpha^{\star}+1}{3}}}=T^{1-(\frac{3}{4}-\epsilon)\frac{\alpha^{\star}+1}{3}}=T^{\frac{3}{4}+\kappa_{1}},$
where $\kappa_{1}=(\alpha^{\star}+1)\epsilon/3-\alpha^{\star}/4>0$.
D.2 When $M=3$
Under $M=3$, the theoretical optimal rate is $T^{(1-\gamma(\alpha^{\star}))/(1-\gamma(\alpha^{\star})^{3})}=T^{9/((\alpha^{\star})^{2}+5\alpha^{\star}+13)}$.
Case of $t_{1}=T^{9/19+\epsilon}$.
Take $\alpha^{\star}=1$. During the first batch, the learner can do no better than pull an arm uniformly at random. We can use the instance $(f_{1}(x)=1,f_{2}(x)=0)$ so that
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\gtrsim t_{1}=T^{\frac{9}{19}+\kappa_{1}},$
where $\kappa_{1}=\epsilon$.
Case of $t_{1}=T^{9/13-\epsilon}$.
Take $\alpha^{\star}=o(1)$. Denote $T_{2}=T^{1/(\gamma(\alpha^{\star})+1)}t_{1}^{\gamma(\alpha^{\star})/(\gamma(\alpha^{\star})+1)}$. Define the events $E_{2}=\{T_{2}<t_{2}\}$ and $E_{3}=\{t_{2}\leq T_{2}\}$. Recall $Q_{i}(\cdot)=\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)\mathbb{P}_{\pi,i,\omega}(\cdot)$ from (13) given $z_{i}>0$ and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. Take $z_{2}=\lceil(36t_{1}2^{2})^{1/(2\beta+d)}\rceil$ and $z_{3}=\lceil(36T_{2}2^{2})^{1/(2\beta+d)}\rceil$. Since $E_{2}$ can be determined by observations up to $t_{1}$, we have
$|Q_{2}(E_{2})-Q_{3}(E_{2})|=|Q_{2}^{t_{1}}(E_{2})-Q_{3}^{t_{1}}(E_{2})|\leq\mathrm{TV}(Q_{2}^{t_{1}},Q_{3}^{t_{1}})\overset{\mathrm{(i)}}{\leq}\frac{1}{2}\sqrt{t_{1}z_{2}^{-(2\beta+d)}}\overset{\mathrm{(ii)}}{\leq}\frac{1}{4},$ (21)
where step (i) applies Lemma 2, and step (ii) is due to the definition of $z_{2}$. Consequently,
$Q_{2}(E_{2})+Q_{3}(E_{3})$ $=Q_{2}(E_{2})-Q_{3}(E_{2})+Q_{3}(E_{2})+Q_{3}(E_{3})$
$\geq-\frac{1}{4}+Q_{3}(E_{2})+Q_{3}(E_{3})=\frac{3}{4},$
where the second step uses inequality (21) and the last equality is because $E_{2}$ and $E_{3}$ form a whole partition of the probability space. Then we would have at least one of $Q_{2}(E_{2})\geq 1/4$ or $Q_{3}(E_{3})\geq 1/4$. If $Q_{2}(E_{2})\geq 1/4$, by Lemma 4 we obtain
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\geq\sup_{f\in\mathcal{C}_{z_{2}}}R_{T_{2}}(\pi;f)\apprge T_{2}z_{2}^{-\beta(1+\alpha^{\star})}\asymp T^{\frac{9}{13}+\kappa_{1}},$
for some $\kappa_{1}>0$. If $Q_{3}(E_{3})\geq 1/4$, we similarly have
$\sup_{f\in\mathcal{F}(\alpha^{\star},\beta)}\mathbb{E}[R_{T}(\pi)]\geq\sup_{f\in\mathcal{C}_{z_{3}}}R_{T_{3}}(\pi;f)\apprge Tz_{3}^{-\beta(1+\alpha^{\star})}\asymp T^{\frac{9}{13}+\kappa_{1}},$
for some $\kappa_{1}>0$.